OpenStudy (anonymous):

joseph and kevin have a combined weight of 120 KG. they then both go on diets. Kevin gains 10% of his previous body weight, and joseph loses 10%. Their combined weight is now 125KG. What are their weights befoe and after?

OpenStudy (lgbasallote):

let x= joseph y = kevin x + y = 120 (y + 0.10y )+(x - 0.10x) = 125 can you do it from here now?

OpenStudy (anonymous):

say, their weights r x kg and y kg respectively so, x+y=120 after the diet course, x will increase by 10%, y will decrease by 10% so, finally, 11/10 x+ 9/10 y= 125 (given) solve these two equations

OpenStudy (anonymous):

I already know what it is, I'm just putting it up here so you guys can relax your brains a bit....

OpenStudy (anonymous):

it is 35/85 before, and 31.5/93.5 after

OpenStudy (anonymous):

x+y=120....1 11/10 x+ 9/10 y=125 11x+9y=1250.....2 from 1, 11x+11y=1320.....3 so, 2y=70, y=35 x=85.. and..

OpenStudy (anonymous):

great.. next time, post a better question :)